OpenStudy (anonymous):

If f(x)=1/x+2 What is, f(3+h)

zepdrix (zepdrix):

\[\Large\rm f(\color{orangered}{x})=\frac{1}{\color{orangered}{x}+2}\]Then,\[\Large\rm f(\color{orangered}{3+h})=\frac{1}{\color{orangered}{3+h}+2}\]

zepdrix (zepdrix):

Or was the function supposed to be \(\Large\rm \frac{1}{x}+2\) ? Was hard to tell with the way you wrote it.

OpenStudy (anonymous):

No the way you wrote it first